OpenStudy (anonymous):

If an object is dropped from a height of 144 ft, the function h(t)=16t^2=144 gives the height of the object after t seconds. When will the object hit the ground?

OpenStudy (anonymous):

16t^2 = 144 t^2 = 144/16 t = 12/4 = 3s

OpenStudy (anonymous):

ok thank you can you explain to me how to do it?

OpenStudy (anonymous):

h(t)=16t^2=144

OpenStudy (anonymous):

using this

OpenStudy (anonymous):

just find t

OpenStudy (anonymous):

oh ok thanks :)

OpenStudy (anonymous):

yw..))